OpenStudy (anonymous):

establish an equation to determine revenue and total cost from a fixed cost of $55, sell price of $15 and labor cost $4. Interpret your result from each.

OpenStudy (anonymous):

55=15x-4

ganeshie8 (ganeshie8):

i dont think fixed cost comes in revenue equation i guess revenue is just R = 15x

ganeshie8 (ganeshie8):

i may be wrong !

OpenStudy (anonymous):

Is that correct?

OpenStudy (anonymous):

revenue equals earned income right?

ganeshie8 (ganeshie8):

revenue is just how much money you make by selling stuff to know profit, you subtract all your expenses from the revenue earned.

OpenStudy (anonymous):

ok, I have a fixed cost of $55 and a selling price of $15 with a labor cost of $4. Could you help me an equation to establish revenue with these figures?

ganeshie8 (ganeshie8):

you're selling each item at a price of $15

ganeshie8 (ganeshie8):

forget about costs, they dont come in Revenue equation

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

so... if you sell "x" items, what revenue you will make (how much money you will get ) ?

OpenStudy (anonymous):

15 times x =??

ganeshie8 (ganeshie8):

thats right ! just label the revenue as R, R = 15x

OpenStudy (anonymous):

oh ok.

ganeshie8 (ganeshie8):

thats the revenue equation

OpenStudy (anonymous):

really? thats it? ok how would we establish an equation for total cost?

ganeshie8 (ganeshie8):

yep thats it !

ganeshie8 (ganeshie8):

total cost = fixed cost + variable cost

ganeshie8 (ganeshie8):

fixed cost you knew already

ganeshie8 (ganeshie8):

variable cost = $4 per EACH item so if you make "x" items, how much the cost will be ?

ganeshie8 (ganeshie8):

just compute it, and put it in an equation. you just need to know these definitions, thats all, then its just simple algebra :D

OpenStudy (anonymous):

55+15x?

ganeshie8 (ganeshie8):

nope. 15 is the selling price. its not labor cost.

OpenStudy (anonymous):

55=y+15x??

ganeshie8 (ganeshie8):

its, total cost C = fixed cost + variable cost C = 55 + 4*x C = 55 + 4x

OpenStudy (anonymous):

what is the formula for profit?

ganeshie8 (ganeshie8):

profit is just, "money you make by selling product" minus "cost in making the product"

ganeshie8 (ganeshie8):

profit = revenue - cost